OpenStudy (anonymous):

if f(x)=x^3 + 1/x and g(x)=f^-1 (x) find the slope of the curve g(x) = f^-1(x) at the point (2,1).

OpenStudy (ranga):

\[Is~f(x) = \frac{ x^3+1 }{ x }~~or~~f(x) = x^3 + \frac{ 1 }{ x }~~?\]

OpenStudy (anonymous):

I didn't even see the \(/x\).

OpenStudy (anonymous):

\[x^3 + \frac{ 1 }{ x }\]

OpenStudy (ranga):

They want you to find g'(x) at the point (2,1). If (2,1) is a point on the inverse function g(x), then (1,2) is a point on the original function f(x). g'(x) at (2,1) = 1 / { f'(x) at (1,2) }

OpenStudy (anonymous):

d/dx (f^-1(a)) = 1/ f' (f^-1(a) ) g(2) = 1 implies f(1) = 2 so d/dx (f^-1(1) ) = 1/ f' (f^-1(1) ) = 1 / f'(2) f'(x) = 3x^2 - 1/x^2 f'(2) = 3(2)^2 - 1/(2)^2 = 47/4 d/dx (f^-1(1)) = 1 / (47/4)) = 4/47

OpenStudy (anonymous):

y u do dis sourwing

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

options are 1, 1/2, 1/3 , -1/2

OpenStudy (anonymous):

oops i evaluate at the wrong point d/dx (f^-1(2)) = 1 / f' (f'^-1)(2) = 1/ f'(1) f'(1) = 2 so 1/2